• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Deur

    Log entry time 16:01:19 on January14,2009

    Entry number 256089

    keyword=Day shift summary

    8:25. DAQ crash. Deadtime is 100%. Rocs rebooted twice but it did not help.
    8:43. DAQ expert called
    9:01 MCC called saying that injector beam study will start soon.
    9:10 DAQ fixed. It seems the problem came from the RICH readout. It was removed from the DAQ and things are fine now. New run started (run 5135). Happex run started as well (run 30628).
    9:14 Run 5135 over (max event # resets by mistake at 100K).
    9:15 started run 5136.
    9:20 Big bite drift chamber trip. Stopped run 5136.
    9:23 A. Camsonne in to try to put back Rich inside DAQ,
    9:28 Started run 5139 (previous runs are tests). Rich still out of DAQ.
    9:29 beam line HV crate tripped. power cycling fixed it. The alarm handler was off and is now restarted (so it's not clear when the crate tripped.
    9:38. Run stopped because we have no beam and A. Camsonne want to finish fix the DAQ.
    run 5140 junk (no beam)
    09:46 pedestal run done.
    runs 514105142 (junk tried to do a cosmic run but with too high scaller values)
    run 5143 cosmic run.
    11:00 Still no beam. We move the target to optic target in order to be ready to check the beam position whenever the beam is back.
    11:03 finished cosmic run 5143, start cosmic run 5144.
    11:25. Still no beam. Called the RC to let him know. RC called PD who said that some injector issues were discovered so the beam study was extended to noon (it would have been nice if MCC had pass on the info to us...).
    11:40 cosmic run 5144 stopped.
    11:50 saw some beam in the Hall (tune beam?). Beam off again after 3 spikes (up to 4 uA).
    12:05, New beam spikes.
    12:09 pulsed beam.
    12:17 14 uA of beam but no word from MCC yet.
    12:21 called MCC to ask the beam status: they have to tune the beam for Hall B first. MCC called back immediately saying they were mistaken and they can actually deliver beam to us now. BPM positions looks good. Asked for 13.5 uA
    12:27 Beam ready for production,
    12:29 Start production run 5146.
    1:25 Big Bite HV trip. Stop run. stop and start BB HV.
    1:29 start run 5147
    1:56 Big Bite HV trip. Stop run. stop and start BB HV.
    1:59 start run 5148
    14:45 Sirish has been working on tuning the beam to reduce the (too high) Compton rates. The beam was quickly tuned (not invasively) and the Compton rates are now good. Part of the background came from injector leakage so Sirish instructed us to run with the injector slits at 28. MCC called and confirmed we are now using this opening for the slits.
    15:05 stop run 5148
    15:06 start run 5149
    15:22 Big bite HV trip. (This time we did not stop and start a new run) All HV back at 15:26.
    15:30 Request Controlled access to MCC to fix the rich and the lumi.